OpenStudy (anonymous):

1) Solve x + 2y = 50 -2x + 3y = 40 5x + 5y + 5z = 300 2) Solve x + y + z = -6 -2x – 2y – 2z = 12 5x + 5y + 5z = -30

OpenStudy (hba):

Well,Well,Well. This question is quite easy. We should use eliminaton for this :))

OpenStudy (anonymous):

Okay, elimination I can do. The three systems threw me off.

OpenStudy (theviper):

yes Elimination is the key to this question :)

OpenStudy (hba):

Don't worry about that, Multiply eq(1) by 2 and add eq(1) and (2) x + 2y = 50 -2x + 3y = 40 5x + 5y + 5z = 300